OpenStudy (anonymous):

Simplify root(75x^2y^-4) a. (5root3)x/y^2 or b. (5root3)|x|/y^2

jhonyy9 (jhonyy9):

so first ,,root" there mean squarroot ?

jhonyy9 (jhonyy9):

hope that yes so than will try it step by step how you can rewriting 75 using square of a number ?

jhonyy9 (jhonyy9):

you dont like cooperate anything ?

jhonyy9 (jhonyy9):

75 = 3*5^2 ok ?

jhonyy9 (jhonyy9):

75x^2y^-4 do you can rewrite it using just positive exponent ?

jhonyy9 (jhonyy9):

so y^(-4) = 1/(y^4) yes ?

OpenStudy (anonymous):

Oh my, so sorry, I went out! I'm back now though. I solved it out: \[\sqrt{75x^{2}y^{-4}}\] \[\sqrt{75} = 5\sqrt{3} \] \[\sqrt{x^2} = x\] \[\sqrt{y^-4} = \frac{ 1 }{ y^2 }\]

jhonyy9 (jhonyy9):

ok yw

OpenStudy (anonymous):

But this would be: \[\frac{ 5 \sqrt{3}x }{ y^2 }\] Is it possible for it to be: \[\frac{ 5 \sqrt{3}|x| }{ y^2 }\]

jhonyy9 (jhonyy9):

a. is right because root(75x^2y^-4) = sqrt((3*5^2 *x^2)/y^4) = (5xsqrt3)/y^2

OpenStudy (anonymous):

Thank you (:

jhonyy9 (jhonyy9):

yw hope you understand it now sure right

OpenStudy (anonymous):

Yep!

jhonyy9 (jhonyy9):

good luck bye